OpenStudy (raffle_snaffle):

The dipole moment (μ) of HBr (a polar covalent molecule) is 0.797D (debye), and its percent ionic character is 11.8 % . Estimate the bond length of the H−Br bond in picometers. Note that 1 D=3.34×10−30 C⋅m and in a bond with 100% ionic character, Q=1.6×10−19 C.

OpenStudy (raffle_snaffle):

@Photon336

OpenStudy (raffle_snaffle):

@Photon336

OpenStudy (photon336):

We will start with this \[\mu = Q*r \]

OpenStudy (photon336):

\[\frac{ 0.79 D }{ x } = \frac{ 11 }{ 100 }\] they've already given us the %ionic character which means that we can find easily find the how many Debyes it would be if the molecule had 100% ionic character by setting up a proportion which is 7.18D \[7.18Dybe \] 0.79/7.18 = 11% ionic character. \[Q*r = \mu \] \[\frac{ 1 D}{ 3.34x10^{-30} C*m } *(1.60x10^{-19})C*r = 7.18D\] I feel that the question gave us the charge, when we had 100% ionic character, so that's why I did it this way, using the charge for that, because we don't know what the charge was when it's 0.79 Dybe's. here i feel that we are estimating the length by assuming 100% ionic character, and then figuring out what the length is. \[\frac{ (7.18D)(3.34x10^{-30}) C*m }{ (1.60x10^{-19})C*D } = 1.50x10^{-10} m\] \[\frac{ 1.50x10^{-10}m }{ 1x10^{-10} m} = 1.5 Angstroms\]

OpenStudy (photon336):

To confirm and justify my answer these were some bond lengths i found as well as the dipole moment of HBR, in Dybes "HBr 1.41 bond length 0.82 dipole moment"

OpenStudy (raffle_snaffle):

@Photon336

OpenStudy (raffle_snaffle):

For some reason the answer is wrong. Is that to the correct sig figs?

OpenStudy (photon336):

Yeah, I didn't take significant figures into account

OpenStudy (photon336):

I'm guessing the answer has 3 significant figures.

OpenStudy (photon336):

@aaronq advice?

OpenStudy (raffle_snaffle):

Hey thanks for your help.

OpenStudy (raffle_snaffle):

oops I was suppose to convert to picometers haha my bad... That is what it asks for.